OpenStudy (anonymous):

If sin Θ = negative square root 3 over 2 and π < Θ < 3 pi over 2, what are the values of cos Θ and tan Θ? cos Θ = negative 1 over 2; tan Θ = square root 3 cos Θ = negative 1 over 2; tan Θ = −1 cos Θ = square root 3 over 4; tan Θ = −2 cos Θ = 1 over 2; tan Θ = square root 3

OpenStudy (anonymous):

help @amistre64 @kirbykirby

OpenStudy (anonymous):

@mathhelppleease

OpenStudy (anonymous):

@andy8150

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

You can try the property: \(\sin^2\theta+\cos^2\theta = 1\) and then isolate to find \(\cos^2 \theta\) But you will find two solutions for \(\cos \theta\). But you use the information for the restriction on \(\pi\) to determine which one you are talking about. Then, you can find \[ \tan\theta = \frac{\sin \theta}{\cos \theta}\]

OpenStudy (anonymous):

so what would the answer be

OpenStudy (kirbykirby):

\[ \sin^2\theta+\cos^2\theta = 1\\ \cos^2\theta= 1 - \sin^2\theta \\ \cos \theta=\pm \sqrt{1-\sin^2\theta}\]

OpenStudy (kirbykirby):

\[ \cos \theta = \pm \sqrt{1-\left( \frac{-\sqrt{3}}{2} \right)^2}\]

OpenStudy (anonymous):

Is it B

OpenStudy (anonymous):

Or D

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

Is It A, B, C or D

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

did you try it?

OpenStudy (anonymous):

yea and i got C

OpenStudy (kirbykirby):

not quite

OpenStudy (anonymous):

ok ill try again

OpenStudy (anonymous):

Is it D

OpenStudy (campbell_st):

just draw a picture